### v1

# Spin-Depairing-Induced Exceptional Fermionic Superfluidity

Abstract
We investigate the non-Hermitian (NH) attractive Hubbard model with spin depairing, which is a spin-resolved asymmetric hopping that nonreciprocally operates spins in the opposite direction. We find that spin depairing stabilizes a superfluid state unique to the NH system. This phase is characterized not only by a finite order parameter, but also by the emergence of exceptional points (EPs) in the momentum space – a feature that starkly contrasts with previously discussed NH fermionic superfluidity, where EPs are absent within the superfluid state and emerge only at the onset of the superfluid breakdown. We uncover the rich mechanism underlying this “exceptional fermionic superfluidity” by analyzing the interplay between EPs and the effective density of states of the complex energy dispersion. Furthermore, we reveal that the exceptional superfluid state breaks down induced by strong spin depairing on the cubic lattice, while it remains robust on the square lattice.
Introduction–Open quantum systems have attracted great interest because they give rise to novel phenomena that are inaccessible in equilibrium settings [1, 2, 3, 4]. In particular, experimental advances in ultracold atoms have realized drastic dissipative many-body phenomena [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], such as loss-induced superfluid-to-Mott insulator transitions [27], anomalous decoherence in many-body systems [28], and dissipation-induced dynamical crossover of the magnetization [29]. Recently, non-Hermitian (NH) systems have gained significant attention as a framework to describe many-body dynamics of open quantum systems [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. One of the most fundamental dissipation effects is the asymmetric hopping [45, 46, 47, 48, 49], which can induce an extreme sensitivity of eigenstates to boundary conditions, called NH skin effects [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. The asymmetric hopping can significantly alter many-body properties [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] as seen in NH skin effects in interacting Hatano-Nelson models [86, 87, 88, 89, 90, 91, 92, 93, 94], breakdown of the Mott insulator [95], and spin-depairing-induced phase transitions [96, 97, 98].
In NH systems, exceptional points (EPs), which are the unique nonequilibrium singularities where the eigenvalues and eigenstates of the NH Hamiltonian coalesce, can ubiquitously emerge in various physical setups [99, 100, 101, 48, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111]; the laser intensity is extremely enhanced at EPs [112, 113], anomalous bulk Fermi arcs can emerge due to the lifetime effect of electrons [114, 115, 116, 117], and unique NH topological phenomena can occur due to the symmetry protection [118, 119, 120, 121, 122]. One of the salient features in NH many-body phenomena is the nonequilibrium phase transitions accompanied by EPs [123, 124, 125, 93], as seen in, e.g., parity-time symmetric systems [126, 86], Kondo problems [127], Bose-Einstein condensates [128, 129], and Mott insulators [130]. In particular, NH fermionic superfluidity (NH-SF) has attracted considerable attention as it exhibits reentrant superfluid phase transitions associated with EPs protected by an emergent NH symmetry induced by the continuous quantum Zeno effect [131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142]. However, in NH-SF, EPs only emerge at the phase boundary, where the metastable superfluid state breaks down [131], the fact of which makes it difficult to experimentally access EPs on top of NH-SF.
Figure 1: Phase diagrams of the attractive Hubbard model with spin depairing on (a) cubic and (b) square lattices. $\gamma$ and $U$ stand for the rate of spin depairing and the strength of attraction, respectively. Spin depairing induces a stable exceptional superfluid state, which is not realized in the Hubbard model with two-body dissipation discussed in previous studies [131]. The corresponding phase diagrams for the model with two-body dissipation $\gamma_{2}$ on the cubic and square lattices are shown in (c) and (d), respectively, for comparison.
In this Letter, we study the NH attractive Hubbard model to find that spin depairing stabilizes the superfluid state unique to the NH system, which we refer to as exceptional fermionic superfluidity. This phase is characterized not only by a finite order parameter, but also by the emergence of EPs in the momentum space. This is in stark contrast to the conventional NH-SF (see Fig. 1). To clarify the origin of the exceptional superfluid state, we employ the NH-BCS theory, introducing the effective density of states (DOS) in the complex energy plane. We find that the interplay between EPs and the effective DOS gives rise to the rich behavior of the exceptional SF. Furthermore, we demonstrate that the exceptional superfluid state breaks down induced by strong spin depairing on the cubic lattice, while it remains robust on the square lattice.
Figure 2: Spin-resolved asymmetric hopping (spin depairing) given in Eq. (1). The hopping amplitude of the fermion with an up spin along the positive (negative) $x,y,z$ directions is described by $-t(1+\gamma)$ [$-t(1-\gamma)$]. In contrast, the hopping amplitude for down-spin fermions is the opposite of that for the up-spin fermions.
NH-BCS theory with spin depairing–We consider the attractive Hubbard model with spin-resolved Hatano-Nelson-type asymmetric hopping [45, 46, 47, 96, 97, 98] as
$H=$ $-t\sum_{\langle i,j\rangle,\sigma}\left[(1+\sigma\gamma)c_{i\sigma}^{\dagger}c_{j\sigma}+(1-\sigma\gamma)c_{j\sigma}^{\dagger}c_{i\sigma}\right]$
$-U\sum_{i}\left(n_{i\uparrow}-\frac{1}{2}\right)\left(n_{i\downarrow}-\frac{1}{2}\right),$ (1)
where $c_{i\sigma}^{\dagger}(c_{i\sigma})$ is the creation (annhilation) operator of a fermion with spin $\sigma=\uparrow,\downarrow$ at site $i$ and $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}$ is the particle number operator. The summation is taken over nearest-neighbor site pair $\langle i,j\rangle$, $t$ is the hopping amplitude, $U\>(>0)$ is the interaction strength, and $\gamma\;(0\leq\gamma\leq 1)$ is the rate of spin depairing. The spin-resolved asymmetric hopping is schematically shown in Fig. 2. Such effective NH descriptions have been employed in the context of dissipative ultracold atoms on an optical lattice [143, 2], and the asymmetric hopping can be implemented by the gauge flux [45, 46, 47, 95, 96] or the collective one-body loss [48, 142, 98].
By performing the Fourier transformation, the kinetic term of Eq. (1) is rewritten as
$H^{\mathrm{kin}}=\sum_{\bm{k}}\left(\epsilon_{\bm{k}}c_{\bm{k}\uparrow}^{\dagger}c_{\bm{k}\uparrow}+\epsilon_{\bm{k}}^{\ast}c_{\bm{k}\downarrow}^{\dagger}c_{\bm{k}\downarrow}\right),$ (2)
where the energy dispersion is given as
$\epsilon_{\bm{k}}=-2t\sum_{\alpha=1}^{d}\left(\cos k_{\alpha}+i\gamma\sin k_{\alpha}\right),$ (3)
where $d$ is the dimension. Note that the presence of spin depairing makes the energy dispersion complex and spin-dependent. Throughout this paper, we set $t$ as the unit of energy.
To analyze the ground state for the NH system, we employ the NH-BCS theory [131]. We start with a path-integral representation of the partition function as
$Z=\int\mathcal{D}\bar{c}\mathcal{D}c\exp(-S),$ (4)
$S=\int_{0}^{\beta^{\prime}}\mathrm{d}\tau\sum_{i}[\bar{c}_{i\sigma}\partial_{\tau}c_{i\sigma}+H(\bar{c}_{i\sigma},c_{i\sigma})],$ (5)
where $\beta^{\prime}$ is a parameter characteristic of the statistical weight of eigenstates, and $H(\bar{c}_{i\sigma},c_{i\sigma})$ is obtained by replacing the fermionic operators with Grassman variables. We then use the Hubbard-Stratonovich transformation to introduce auxiliary fields $\Delta$ and $\bar{\Delta}$. Integrating out the fermionic degrees of freeedom and taking the saddle point of the action with respect to $\Delta$ and $\bar{\Delta}$, we obtain the NH gap equation as
$\frac{1}{U}=\frac{1}{N}\sum_{\bm{k}}\frac{1}{2E_{\bm{k}}},$ (6)
where $N$ is the number of sites, $E_{\bm{k}}=\sqrt{\epsilon_{\bm{k}}^{2}+\Delta\bar{\Delta}}$, and we have taken the limit $\beta^{\prime}\to\infty$. Note that the gauges of the order parameters $\Delta$ and $\bar{\Delta}$ can be taken to be real. This is in contrast to the result for the conventional NH-SF induced by two-body dissipation, where the order parameters are intrinsically complex [131]. In the following, we set $\Delta_{0}=\Delta=\bar{\Delta}\in\mathbb{R}$.
By using the order parameter $\Delta_{0}$ that satisfies the NH gap equation (6), the NH-BCS Hamiltonian can be rewritten as
$H^{\mathrm{BCS}}=\sum_{\bm{k}}\begin{pmatrix}c_{\bm{k}\uparrow}^{\dagger}&c_{-\bm{k}\downarrow}\end{pmatrix}\begin{pmatrix}\epsilon_{\bm{k}}&\Delta_{0}\\ \Delta_{0}&-\epsilon_{\bm{k}}\end{pmatrix}\begin{pmatrix}c_{\bm{k}\uparrow}\\ c_{-\bm{k}\downarrow}^{\dagger}\end{pmatrix}.$ (7)
The ground state energy is given as
$E_{\mathrm{SF}}=\frac{\Delta_{0}^{2}}{U}-\sum_{k}[E_{\bm{k}}-\epsilon_{\bm{k}}].$ (8)
The condensation energy is given by $E_{\mathrm{cond}}=E_{\mathrm{SF}}-E_{\mathrm{N}}$, where $E_{\mathrm{N}}$ is the energy for the normal state with $\Delta_{0}=0$. Here, we remark that we obtain the real condensation energy that gives the same decay rate between the superfluid and normal states. This is distinct from that for the conventional NH system, where the superfluid state usually decays faster than the normal state. Note that the effective Hamiltonian (7) becomes non-diagonalizable at $E_{\bm{k}}=0$, indicating the presence of the EPs with $\epsilon_{k}=\pm i\Delta_{0}$.
Figure 3: (a) [(c)] Contour plot of the effective DOS (9) on the cubic (square) lattice for $\gamma=1$. Blue region indicates the effective DOS being zero. (b)[(d)] Cross section of the effective DOS.
To solve the gap equation (6), it is convenient to introduce the effective DOS of a complex variable $\epsilon$ as
$D_{\gamma}(\epsilon)$ $=\frac{1}{N}\sum_{\bm{k}}\delta\left({\rm Re}\,\epsilon-{\rm Re}\,\epsilon_{\bm{k}}\right)\delta({\rm Im}\,\epsilon-{\rm Im}\,\epsilon_{\bm{k}}),$ (9)
where $\delta(x)$ is the delta function. The effective DOS satisfies $D_{\gamma}(\epsilon)=D_{\gamma=1}({\rm Re}\,\epsilon+i{\rm Im}\,\epsilon/\gamma)/\gamma$ and $D_{\gamma}(\epsilon)=D_{\gamma}(\epsilon^{*})$, and it is finite at the elliptic region in the complex-$\epsilon$ plane. Some details are provided in Appendix A. Figure 3 shows the effective DOS in the systems with $\gamma=1$ on the cubic and square lattices. We find that both lattices have the isotropic effective DOS in the complex-$\epsilon$ plane, while distinct singularities emerge between them. The effective DOS logarithmically diverges at $|\epsilon|=2$ in the cubic lattice case, while algebraically diverges at $|\epsilon|=0$ and $|\epsilon|=4$ in the square lattice case, which is clearly found in the cross section of the effective DOS shown in Figs. 3(b) and (d), respectively. The qualitative difference in the effective DOS affects the ground state phase diagram, which will be discussed later. The gap equation is then given as
$\frac{1}{U}=\int\!\!\!\!\int\mathrm{d}\text{Re}\,\epsilon\>\mathrm{d}\text{Im}\,\epsilon\frac{D_{\gamma}(\epsilon)}{2\sqrt{\epsilon^{2}+\Delta_{0}^{2}}}.$ (10)
Figure 4: (a) Normalized order parameter and (b) condensation energy for the NH system with the constant effective DOS. Exceptional SF is realized in the green shaded area.
Analytical results for the spin-depairing-induced exceptional SF–To understand how the attractive interaction induces the SF unique to the NH system with spin depairing, we first consider the effective DOS in the absence of singularities. Specifically, we introduce an effective DOS defined in the complex-$\epsilon$ plane, which takes a constant value within a circular domain:
$D(\epsilon)=\begin{cases}1/(\pi\epsilon_{R}^{2}),\;&\;(|\epsilon|\leq\epsilon_{R}),\\ 0,\;&\;(|\epsilon|>\epsilon_{R}),\end{cases}$ (11)
where the cutoff $\epsilon_{R}$ has been introduced so that the normalization condition $\int\!\!\!\int\mathrm{d}\text{Re}\epsilon\mathrm{d}\text{Im}\epsilon D(\epsilon)=1$ is satisfied. Integrating over $\epsilon$ in the NH gap equation (6), we obtain the following equation
$\frac{1}{U^{\prime}}=\begin{cases}\frac{\pi}{2\Delta^{\prime}},\;&\;(\Delta^{\prime}>1),\\ \frac{\pi}{2\Delta^{\prime}}+\sqrt{1-\Delta^{\prime 2}}-\frac{1}{\Delta^{\prime}}\tan^{-1}(\frac{\sqrt{1-\Delta^{\prime 2}}}{\Delta^{\prime}}),\;&\;(\Delta^{\prime}\leq 1),\end{cases}$ (12)
where we have introduced the parameters $U^{\prime}\equiv U/(\pi\epsilon_{R})$ and $\Delta^{\prime}\equiv\Delta_{0}/\epsilon_{R}$, for simplicity. We obtain the condensation energy as
[…]
(for the detailed derivation, see Appendix B).
The results for this simple case are shown in Fig. 4. When $U^{\prime}$ is large, the strong attraction stabilizes the trivial SF state with finite $\Delta^{\prime}$. This is consistent with the fact that the condensation energy is negative. In this SF phase, $\Delta_{0}>\epsilon_{R}$ and the quasiparticle energy $E_{k}$ remains finite in the complex energy plane. As the attraction decreases, the order parameter $\Delta_{0}$ also decreases and eventually reaches $\Delta_{0}=\epsilon_{R}$ at $U^{\prime}=2/\pi\sim 0.637$. This means that the EPs emerge at the edge of the circular DOS, i.e., $\epsilon=\pm i\Delta_{0}$, which is nothing but the emergence of the exceptional SF. Remarkably, this phase has not been discussed in nonequilibrium systems, to the best of our knowledge. As the attraction decreases further, the order parameter continues to decrease and the EPs move toward the origin in the complex-$\epsilon$ plane. Finally, $\Delta_{0}$ vanishes at the critical interaction strength $U^{\prime}=U_{c}^{\prime}\equiv 1/2$, as shown in Fig. 4(a). Near this point, the order parameter exhibits mean-field asymptotic behavior $\Delta^{\prime}\sim\sqrt{(U^{\prime}-U_{c}^{\prime})}$. The key result here is that the exceptional SF – characterized by the finite order parameter and the presence of EPs inside the complex-$\epsilon$ plane – is indeed stabilized within a finite interaction range. In the following, we examine whether such a exceptional SF also arises in the systems with cubic and square lattices and discuss the effects of singularities in the effective DOS.
Figure 5: Numerical results for the NH system on the cubic lattice with $\gamma=0.5$. (a) Order parameter $\Delta_{0}$ as a function of the attraction $U$. Exceptional SF is realized in the green shaded area, where EPs appear at $\epsilon=\pm i\Delta_{0}$. (b) Contour plot of the effective DOS (9) and (c) its cross section at $\mathrm{Re}\epsilon=0$. The marks in (b) and (c) stand for the points $\mathrm{Im}\epsilon=i\Delta_{0}$ at $\mathrm{Re}\epsilon=0$ by using the corresponding value of $\Delta_{0}$ shown in (a). (d) Exceptional lines in the momentum space when $U=5.57$ and $7.02$.
Exceptional SF on the cubic lattice–We first consider the NH system on the cubic lattice. Figure 5(a) shows the order parameter $\Delta_{0}$ as a function of the attractive interaction for the system with $\gamma=0.5$. When $U=8$, a stable SF state is realized with large $\Delta_{0}\sim 3.53$. In this case, the position $\epsilon=\pm i\Delta_{0}$ is located outside the elliptic region with finite $D(\epsilon)$ in the complex-$\epsilon$ plane, as shown in Figs. 5(a)-(c) (see the blue circle). This means the absence of EPs in the superfluid state. As the attraction decreases, the order parameter $\Delta_{0}$ also decreases, and eventually reaches $\Delta_{0}=3$ when $U\sim 7.02$. In this case, EPs emerge at $\epsilon=\pm i\Delta_{0}$ in the complex-$\epsilon$ plane (see the pink cross mark), indicating the onset of exceptional SF. The corresponding EPs in the momemtum space emerges at ±$\pm(\pi/2,\pi/2,\pi/2)$, which are shown as the pink circles in Fig. 5(d). When $U\sim 5.57$, the exceptional SF is realized with $\Delta_{0}\sim 2$. The corresponding EPs are observed as the lines in the momentum space in Fig. 5(d). As the attraction further decreases, the order parameter suddenly vanishes at $\Delta_{0}=1$. Since this singularity is distinct from the critical behavior for the system with the constant effective DOS, this behavior can be attributed to the logarithmic divergence of the effective DOS at that energy (see the red star). As a result, the exceptional SF becomes unstable and the normal state is realized instead. This is in contrast to the Hermitian limit where the superfluid state remains stable for any attractive interaction $U$.
By performing similar calculations for several $\gamma$, we obtain the phase diagram, as shown in Fig. 1(a). We find that an exotic exceptional SF, characterized by the finite order parameter and the quasiparticle energy exhibiting nodes at EPs, is stabilized over a finite range of interaction and spin depairing strengths. This exceptional SF is qualitatively distinct from both metastable and stable SF states realized in the NH system with two-body dissipation [131]. In fact, no nodes appear in these states and EPs emerge only when the metastable solution for the superfluid state vanishes, as shown in Fig. 1(c).
Figure 6: Numerical results for the NH system on the square lattice with $\gamma=0.5$. (a) Order parameter $\Delta_{0}$ as a function of the attraction $U$. Exceptional SF is realized in the green shaded area. (b) Contour plot of the effective DOS (9) and (c) its cross section at $\mathrm{Re}\epsilon=0$. The marks in (b) and (c) stand for the points $\mathrm{Im}\epsilon=i\Delta_{0}$ at $\mathrm{Re}\epsilon=0$ by using the corresponding value of $\Delta_{0}$ shown in (a). (d) EPs in the momentum space in the NH system with $U=0,3.95$, and $4.92$.
Exceptional SF on the square lattice–We present the result for the NH system on the square lattice and highlight the qualitative difference from the results for the cubic lattice. In Fig. 6(a), we show the order parameter $\Delta_{0}$ as a function of $U$ for $\gamma=0.5$. We find that the order parameter is finite in the whole region, implying that the spontaneous symmetry breaking always occurs. Furthermore, we find that when $U\leq U_{c}(\sim 4.92)$, the order parameter is less than the edge of the elliptic region with finite DOS i.e. $\Delta_{0}\leq 2$, meaning that the exceptional SF is stabilized. When $U=U_{c}$, EPs emerge at ±$\pm(\pi/2,\pi/2)$ in the momentum space, as shown in Fig. 6(d). Decreasing $U$, each EP split into two EPs and move toward $(0,\pm\pi)$ and (±$(\pm\pi,0)$. It is intriguing that this exceptional SF is stable for the weak coupling regime. This is in contrast to the results for the system on the cubic lattice, where the phase transition occurs between the exceptional SF and the normal state. This is attributed to the structure of the effective DOS. Figures 6(b) and 6(c) show that the effective DOS exhibits two singularities at the origin and the perimeter (see the red star and blue circle). These singularities are characterized by $D(\epsilon)\sim|\epsilon|^{-1}$ for $|\epsilon|\rightarrow+0$ and $D(\epsilon)\sim(2-|\epsilon|)^{-1/2}$ for $|\epsilon|\rightarrow 2-0$. When $\Delta_{0}\sim 2$, the singularity in the effective DOS leads to a cusp in the curve of the order parameter, as shown in Fig. 6(a). In contrast, when $\Delta_{0}\rightarrow 0$, the integral in Eq. (10) tends to diverge due to the strong singularity in the effective DOS at $|\epsilon|=0$. By taking this into account correctly, we find asymptotic behavior $\Delta_{0}\sim\exp(-c_{\gamma}/U)$ with positive constant $c_{\gamma}$ (see Appendix C for details). Therefore, we can say that in the NH system on the square lattice, an infinitesimal attractive interaction induce the exceptional SF, which is clearly found in the phase diagram [see Fig. 1(b)].
Finally, we would like to comment on the exceptional SF state. We have found that the EPs appear in the momentum space as exceptional lines in the cubic lattice, and exceptional points in the square lattice. This indicates that the dimensionality of EPs is given by $d-2$ in a $d$-dimensional system. This result contrasts with that of the conventional NH systems with two-body dissipation, where the dimensionality is typically given by $d-1$ [131]. We also note that gapless excitations at isolated points in momentum space appear in (usually metastable) superfluid state when equilibrium SF is considered under magnetic fields or with mass imbalance [144, 145, 146, 147, 148]. However, we emphasize that spin-depairing-induced exceptional SF is always stable due to the negative condensation energy and is a unique feature of NH systems characterized by EPs.
Conclusion–In this Letter, we have investigated the NH attractive Hubbard model with spin depairing and found the emergence of stable exceptional SF, whose behavior is characterized by the interplay between EPs and the effective DOS. Moreover, we have elucidated that the spin-depairing-induced phase transitions occur on the cubic lattice, while exceptional SF is stable for arbitrary strength of attraction on the square lattice. As we have focused on the analysis of the order parameter, the NH physical quantities in various setups, such as the susceptibility and the superfluid weight, are worth studying to further understand the exceptional SF. Moreover, it is intriguing how the exceptional SF appears for more complex energy-band structures. As the nonreciprocity in Liouvillian many-body dynamics is actively investigated [52, 149, 150, 151, 152, 153, 154, 155], it is also interesting to study nonreciprocal dissipation effects on the superfluid state in nonequilibrium steady states as well as the exceptional SF in general Markovian dynamics [156, 157] that may exhibit Liouvillian EPs [158].
Acknowledgements.
We thank Kohei Kawabata, Norio Kawakami, and Masaya Nakagawa for fruitful discussions. This work was supported by Grant-in-Aid for Scientific Research from JSPS, KAKENHI Grants Nos. JP23K19031, JP25K17327 (K.Y.), and JP22K03525, JP25H01521, JP25H01398 (A.K.). S.T. was supported by the Sasakawa Scientific Research Grant from the Japan Science Society. K.Y. was also supported by Toyota RIKEN Scholar program, Murata Science and Education Foundation, Public Promoting Association Kura Foundation, Hirose Foundation, the Precise Measurement Technology Promotion Foundation, and the Fujikura Foundation.

### v3

@@
 Recently, non-Hermitian (NH) systems have gained significant attention as a framework to describe many-body dynamics of open quantum systems [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
-One of the most fundamental dissipation effects is the asymmetric hopping [45, 46, 47, 48, 49], which can induce an extreme sensitivity of eigenstates to boundary conditions, called NH skin effects [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72].
-The asymmetric hopping can significantly alter many-body properties [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] as seen in NH skin effects in interacting Hatano-Nelson models [86, 87, 88, 89, 90, 91, 92, 93, 94], breakdown of the Mott insulator [95], and spin-depairing-induced phase transitions [96, 97, 98].
-In NH systems, exceptional points (EPs), which are the unique nonequilibrium singularities where the eigenvalues and eigenstates of the NH Hamiltonian coalesce, can ubiquitously emerge in various physical setups [99, 100, 101, 48, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111]; the laser intensity is extremely enhanced at EPs [112, 113], anomalous bulk Fermi arcs can emerge due to the lifetime effect of electrons [114, 115, 116, 117], and unique NH topological phenomena can occur due to the symmetry protection [118, 119, 120, 121, 122].
+One of the most fundamental dissipation effects is the asymmetric hopping [45, 46, 47, 48, 49], which can induce an extreme sensitivity of eigenstates to boundary conditions, called NH skin effects [50, 51].
+The asymmetric hopping can significantly alter many-body properties [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63] as seen in NH skin effects in interacting Hatano-Nelson models [64, 65], breakdown of the Mott insulator [66], and spin-depairing-induced phase transitions [67, 68, 69].
+In NH systems, exceptional points (EPs), which are the unique nonequilibrium singularities where the eigenvalues and eigenstates of the NH Hamiltonian coalesce, can ubiquitously emerge in various physical setups [70, 48, 71, 72, 73]; the laser intensity is extremely enhanced at EPs [74], anomalous bulk Fermi arcs can emerge due to the lifetime effect of electrons [75, 76, 77], and unique NH topological phenomena can occur due to the symmetry protection [78, 79, 80, 81, 82].
 One of the salient features in NH many-body phenomena is the nonequilibrium phase transitions accompanied by EPs [83, 84, 85, 86], as seen in, e.g., parity-time symmetric systems [87, 88], Kondo problems [89], Bose-Einstein condensates [90, 91], and Mott insulators [92].
-In particular, NH fermionic superfluidity (NH-SF) has attracted considerable attention as it exhibits reentrant superfluid phase transitions associated with EPs protected by an emergent NH symmetry induced by the continuous quantum Zeno effect [131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142].
+However, there are few mechanisms in which many-body phases associated with EPs emerge and such a realization of many-body EPs has been a major challenge even in prototypical NH many-body physics.
+In particular, NH fermionic superfluidity (NH-SF) has attracted considerable attention as it exhibits reentrant superfluid phase transitions associated with EPs protected by an emergent NH symmetry induced by the continuous quantum Zeno effect [93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103].
 However, in NH-SF, EPs only emerge at the phase boundary, where the metastable superfluid state breaks down [93], the fact of which makes it difficult to experimentally access EPs on top of NH-SF.
+Since NH fermionic superfluidity can exhibit unconventional yet universal behavior at EPs in fundamental physical quantities, such as in correlation length, compressibility, and correlation functions [93, 99], the mechanism for stable NH fermionic superfluidity associated with EPs can lead to understanding essential physical responses induced by EPs in NH many-body quantum matter.
 Figure 1: Phase diagrams of the attractive Hubbard model with spin depairing on (a) cubic and (b) square lattices.
@@
 Spin depairing induces a stable exceptional superfluid state, which is not realized in the Hubbard model with two-body dissipation discussed in previous studies [93].
-The corresponding phase diagrams for the model with two-body dissipation $\gamma_{2}$ on the cubic and square lattices are shown in (c) and (d), respectively, for comparison.
+The corresponding phase diagrams for the model with two-body dissipation $\gamma_{2}$ on the cubic and square lattices are shown in (c) and (d), respectively [104].
 In this Letter, we study the NH attractive Hubbard model to find that spin depairing stabilizes the superfluid state unique to the NH system, which we refer to as exceptional fermionic superfluidity.
@@
 Furthermore, we demonstrate that the exceptional superfluid state breaks down induced by strong spin depairing on the cubic lattice, while it remains robust on the square lattice.
+The exceptional fermionic superfluidity uncovers a new structure linking EPs and NH fermionic superfluidity, and opens a new avenue towards experimental realization of EPs in NH many-body physics.
 Figure 2: Spin-resolved asymmetric hopping (spin depairing) given in Eq.
 (1).
-The hopping amplitude of the fermion with an up spin along the positive (negative) $x,y,z$ directions is described by $-t(1+\gamma)$ [$-t(1-\gamma)$].
+The hopping amplitude of the fermion with an up spin along the positive (negative) $x,y,z$ directions is described by $t(1+\gamma)$ $t(1-\gamma)$].
 In contrast, the hopping amplitude for down-spin fermions is the opposite of that for the up-spin fermions.
@@
 where $c_{i\sigma}^{\dagger}(c_{i\sigma})$ is the creation (annhilation) operator of a fermion with spin $\sigma=\uparrow,\downarrow$ at site $i$ and $n_{i\sigma}=c_{i\sigma}^{\dagger}c_{i\sigma}$ is the particle number operator.
-The summation is taken over nearest-neighbor site pair $\langle i,j\rangle$, $t$ is the hopping amplitude, $U\>(>0)$ is the interaction strength, and $\gamma\;(0\leq\gamma\leq 1)$ is the rate of spin depairing.
+The summation is taken over nearest-neighbor site pair $\langle i,j\rangle$, $t$ is the hopping amplitude, U(>0)U\>(>0) is the interaction strength, and $\gamma\;(0\leq\gamma\leq 1)$ is the rate of spin depairing.
 The spin-resolved asymmetric hopping is schematically shown in Fig.
 2.
-Such effective NH descriptions have been employed in the context of dissipative ultracold atoms on an optical lattice [143, 2], and the asymmetric hopping can be implemented by the gauge flux [45, 46, 47, 95, 96] or the collective one-body loss [48, 142, 98].
+Such effective NH descriptions have been employed in the context of dissipative ultracold atoms on an optical lattice [105, 2], and the asymmetric hopping can be implemented by the gauge flux [45, 46, 47, 66, 67] or the collective one-body loss [48, 103, 69] (see Appendix A in End Matter for the experimental implementation of our system).
 By performing the Fourier transformation, the kinetic term of Eq. (1) is rewritten as
@@
 By using the order parameter $\Delta_{0}$ that satisfies the NH gap equation (6), the NH-BCS Hamiltonian can be rewritten as
-$H^{\mathrm{BCS}}=\sum_{\bm{k}}\begin{pmatrix}c_{\bm{k}\uparrow}^{\dagger}&c_{-\bm{k}\downarrow}\end{pmatrix}\begin{pmatrix}\epsilon_{\bm{k}}&\Delta_{0}\\ \Delta_{0}&-\epsilon_{\bm{k}}\end{pmatrix}\begin{pmatrix}c_{\bm{k}\uparrow}\\ c_{-\bm{k}\downarrow}^{\dagger}\end{pmatrix}.$ (7)
+$H^{\mathrm{BCS}}=\sum_{\bm{k}}\begin{pmatrix}c_{\bm{k}\uparrow}^{\dagger}&c_{-\bm{k}\downarrow}\end{pmatrix}\begin{pmatrix}\epsilon_{\bm{k}}&\Delta_{0}\\ \Delta_{0}&-\epsilon_{\bm{k}}\end{pmatrix}\begin{pmatrix}c_{\bm{k}\uparrow}\\ c_{-\bm{k}\downarrow}^{\dagger}\end{pmatrix}$.
+(7)
 The ground state energy is given as
@@
 The effective DOS satisfies $D_{\gamma}(\epsilon)=D_{\gamma=1}({\rm Re}\,\epsilon+i{\rm Im}\,\epsilon/\gamma)/\gamma$ and $D_{\gamma}(\epsilon)=D_{\gamma}(\epsilon^{*})$, and it is finite at the elliptic region in the complex-$\epsilon$ plane.
-Some details are provided in Appendix A.
+Some details are provided in Supplemental Material [106].
 Figure 3 shows the effective DOS in the systems with $\gamma=1$ on the cubic and square lattices.
@@
 […]
-(for the detailed derivation, see Appendix B).
+(for the detailed derivation, see Supplemental Material [106]).
 The results for this simple case are shown in Fig.
@@
 The key result here is that the exceptional SF – characterized by the finite order parameter and the presence of EPs inside the complex-$\epsilon$ plane – is indeed stabilized within a finite interaction range.
+Our findings of the exceptional SF reveal a qualitatively new connection between EPs and NH many-body physics, namely, the emergence of the fermionic superfluidity accompanied by EPs within a stable finite parameter region.
+We note that in nonequilibrium systems, a “phase” is commonly defined as one that can be adiabatically connected to the corresponding conventional phase through the gradual introduction of dissipation.
 In the following, we examine whether such a exceptional SF also arises in the systems with cubic and square lattices and discuss the effects of singularities in the effective DOS.
@@
 In this case, EPs emerge at $\epsilon=\pm i\Delta_{0}$ in the complex-$\epsilon$ plane (see the pink cross mark), indicating the onset of exceptional SF.
-The corresponding EPs in the momemtum space emerges at ±$\pm(\pi/2,\pi/2,\pi/2)$, which are shown as the pink circles in Fig.
+The corresponding EPs in the momemtum space emerges at $\pm(\pi/2,\pi/2,\pi/2)$, which are shown as the pink circles in Fig.
 5(d).
@@
 Since this singularity is distinct from the critical behavior for the system with the constant effective DOS, this behavior can be attributed to the logarithmic divergence of the effective DOS at that energy (see the red star).
-As a result, the exceptional SF becomes unstable and the normal state is realized instead.
+As a result, the exceptional SF becomes unstable and the normal state is realized instead (see Appendix B in End Matter).
 This is in contrast to the Hermitian limit where the superfluid state remains stable for any attractive interaction $U$.
@@
 $\Delta_{0}\leq 2$, meaning that the exceptional SF is stabilized.
-When $U=U_{c}$, EPs emerge at ±$\pm(\pi/2,\pi/2)$ in the momentum space, as shown in Fig.
+When $U=U_{c}$, EPs emerge at $\pm(\pi/2,\pi/2)$ in the momentum space, as shown in Fig.
 6(d).
-Decreasing $U$, each EP split into two EPs and move toward $(0,\pm\pi)$ and (±$(\pm\pi,0)$.
+Decreasing $U$, each EP split into two EPs and move toward $(0,\pm\pi)$ and $(\pm\pi,0)$.
 It is intriguing that this exceptional SF is stable for the weak coupling regime.
@@
 (10) tends to diverge due to the strong singularity in the effective DOS at $|\epsilon|=0$.
-By taking this into account correctly, we find asymptotic behavior $\Delta_{0}\sim\exp(-c_{\gamma}/U)$ with positive constant $c_{\gamma}$ (see Appendix C for details).
+By taking this into account correctly, we find asymptotic behavior $\Delta_{0}\sim\exp(-c_{\gamma}/U)$ with positive constant $c_{\gamma}$ (see Supplemental Material [106] for details).
 Therefore, we can say that in the NH system on the square lattice, an infinitesimal attractive interaction induce the exceptional SF, which is clearly found in the phase diagram [see Fig.
 1(b)].
+The intrinsic differences in the phase diagrams between the cubic and square lattices originate from differences in the effective DOS, which arise from the underlying lattice geometry.
 Finally, we would like to comment on the exceptional SF state.
 We have found that the EPs appear in the momentum space as exceptional lines in the cubic lattice, and exceptional points in the square lattice.
-This indicates that the dimensionality of EPs is given by $d-2$ in a $d$-dimensional system.
-This result contrasts with that of the conventional NH systems with two-body dissipation, where the dimensionality is typically given by $d-1$ [131].
+This indicates that the dimensionality of EPs is given by d2d-2 in a $d$-dimensional system.
+This result contrasts with that of the conventional NH systems with two-body dissipation, where the dimensionality is typically given by d1d-1 [93].
 We also note that gapless excitations at isolated points in momentum space appear in (usually metastable) superfluid state when equilibrium SF is considered under magnetic fields or with mass imbalance [107, 108, 109, 110, 111].
@@
 Moreover, we have elucidated that the spin-depairing-induced phase transitions occur on the cubic lattice, while exceptional SF is stable for arbitrary strength of attraction on the square lattice.
+We have shown that the existence of the exceptional SF is established in the thermodynamic limit, where edge effects are excluded [50, 51].
+Therefore, within the present framework, we are not able to provide a complete numerical analysis under open boundary conditions (OBCs).
+To incorporate edge effects, one would need to reconstruct the NH BCS theory under open boundary conditions by introducing site-dependent and inhomogeneous order parameters, which is highly nontrivial.
+We note that, though some works consider NH skin effect under OBCs by assuming bulk order parameters, the order parameters are not self-consistently determined within BCS theory and this approach should be distinguished from our study.
 As we have focused on the analysis of the order parameter, the NH physical quantities in various setups, such as the susceptibility and the superfluid weight, are worth studying to further understand the exceptional SF.
 Moreover, it is intriguing how the exceptional SF appears for more complex energy-band structures.
-As the nonreciprocity in Liouvillian many-body dynamics is actively investigated [52, 149, 150, 151, 152, 153, 154, 155], it is also interesting to study nonreciprocal dissipation effects on the superfluid state in nonequilibrium steady states as well as the exceptional SF in general Markovian dynamics [156, 157] that may exhibit Liouvillian EPs [158].
+Finally, since the nonreciprocity in Liouvillian many-body dynamics is actively investigated [112, 113, 114, 115, 116, 117, 118, 119], it is also interesting to study nonreciprocal dissipation effects on the superfluid state in nonequilibrium steady states as well as the exceptional SF in general Markovian dynamics [120, 121] that may exhibit Liouvillian EPs [122].
 Acknowledgements.
-We thank Kohei Kawabata, Norio Kawakami, and Masaya Nakagawa for fruitful discussions.
+We thank Masaya Nakagawa, Kohei Kawabata, and Norio Kawakami for fruitful discussions.
 This work was supported by Grant-in-Aid for Scientific Research from JSPS, KAKENHI Grants Nos.
-JP23K19031, JP25K17327 (K.Y.), and JP22K03525, JP25H01521, JP25H01398 (A.K.).
+JP25K17327 (K.Y.), and JP22K03525, JP25H01521, JP25H01398 (A.K.).
 S.T.
-was supported by the Sasakawa Scientific Research Grant from the Japan Science Society.
+was supported by the Sasakawa Scientific Research Grant from the Japan Science Society and JST SPRING, Japan Grant Number JPMJSP2180.
 K.Y.
-was also supported by Toyota RIKEN Scholar program, Murata Science and Education Foundation, Public Promoting Association Kura Foundation, Hirose Foundation, the Precise Measurement Technology Promotion Foundation, and the Fujikura Foundation.
+was also supported by Murata Science and Education Foundation, Hirose Foundation, the Precise Measurement Technology Promotion Foundation, and the Fujikura Foundation.
+Appendix B: Discontinuous jump of the order parameter on a cubic lattice
+The gap equation [Eq. (10) in the main text] is rewritten as,
+$U$ $=F(\Delta_{0}),$ (21)
+$F(\Delta_{0})$ $=\left(\int\!\!\!\!\int\frac{D_{\gamma}(\epsilon)}{2\sqrt{\epsilon^{2}+\Delta_{0}^{2}}}\;\mathrm{d}\text{Re}\,\epsilon\>\mathrm{d}\text{Im}\,\epsilon\right)^{-1}.$ (22)
+In Fig.
+8, we show $F(\Delta_{0})$ for the systems on the cubic and square lattices.
+In the large-$\Delta_{0}$ region, $F(\Delta_{0})$ increases monotonically, indicating the existence of a trivial physical solution and hence the realization of the superfluid state.
+In the cubic lattice, $F(\Delta_{0})$ exhibits a minimum, implying that no physical solution exists for $U<U_{c}(U_{c}\sim 4.59)$, which leads to the discontinuous jump in the order parameter $\Delta_{0}$.
+This nonmonotonic behavior arises from the interplay between the singularity of the effective DOS and EPs.
+When $\Delta_{0}>1$, EPs appear at $\epsilon=i\Delta_{0}$ in the complex $\epsilon$ plane.
+This means that the integrand in $F$ diverges at $\epsilon=i\Delta_{0}$, and $D_{\gamma}(\epsilon\sim i\Delta_{0})$ dominates $F$.
+Since the singularity of the effective DOS is located at $\epsilon=i$, this yields the minimum in $F$, as shown in Fig.
+8(a).
+We note that the solution for $\Delta_{0}<1$, which is shown as the dashed line in Fig.
+8, is not physical and should be omitted because they does not adiabatically connect to the solution in the Hermitian limit $\gamma\to 0$.
+In the square lattice case, $F(\Delta_{0})$ approaches zero as $\Delta_{0}\rightarrow 0$, so the superfluid solution always exists although the order parameter is extremely small for small $U$.
+Figure 8: Lines represent $F(\Delta_{0})$, defined in Eq.
+(22), on (a) cubic and (b) square lattices.
+In (a), solid (dashed) line corresponds to the physical (unphysical) solution of the self-consistency equation $U=F(\Delta_{0})$.
+Supplemental Material for
+”Spin-Depairing-Induced Exceptional Fermionic Superfluidity”
+Appendix A Detailed derivation of the effective density of states for the square lattice
+We explain the detailed derivation of the effective density of states (DOS) for the square lattice.
+First, we show the relation between the effective DOS for $\gamma=1$ and that for $\gamma=\gamma_{0}\neq 1$.
+We can rewrite the effective DOS given in Eq.
+(9) in the main text as
+$D_{\gamma}(\epsilon=x+iy)=\frac{1}{N}\sum_{\bm{k}}\delta(x-\text{Re}\epsilon_{\bm{k}})\delta(y-\text{Im}\epsilon_{\bm{k}}).$ (S1)
+Using the relation $\text{Im}\epsilon_{\bm{k}}|_{\gamma=\gamma_{0}}=\gamma_{0}\text{Im}\epsilon_{\bm{k}}|_{\gamma=1}$ and $\delta(ax)=\delta(x)/a$ ($a>0$), we obtain the following relation:
+$D_{\gamma_{0}}(x+iy)=\frac{1}{\gamma_{0}}D_{\gamma=1}(x+iY),$ (S2)
+where $Y=y/\gamma_{0}$. This relation indicates that the effective DOS for $\gamma=1$ contains all the information of $D_{\gamma}(\epsilon)$.
+Next, we show the explicit form of the effective DOS on the square lattice for $\gamma=1$. Since the effective DOS for $\gamma=1$ only depends on the absolute value of the energy $\epsilon$, we get
+$D_{\gamma=1}(\epsilon=re^{i\theta})=D(r)=\frac{1}{2\pi rN}\sum_{\bm{k}}\delta(r-|\epsilon_{\bm{k}}|)=\frac{1}{2\pi r}R(r),$ (S3)
+where $R(r)=\frac{1}{N}\sum_{\bm{k}}\delta(r-|\epsilon_{\bm{k}}|)$. We can rewrite $R(r)$ as
+$R(r)=\int_{-\pi}^{\pi}\mathrm{d}k_{y}\int_{-\pi}^{\pi}\mathrm{d}k_{x}\delta\left(r-4\left|\cos\left(\frac{k_{x}-k_{y}}{2}\right)\right|\right).$ (S4)
+Using the relation $\cos(-x)=\cos(x)$ and introducing $p=k_{x}-k_{y}$, we obtain
+$R(r)=2\int_{-\pi}^{\pi}\mathrm{d}k_{y}\int_{0}^{\pi-k_{y}}\mathrm{d}p\>\delta\left(r-4\left|\cos\frac{p}{2}\right|\right).$ (S5)
+Then, using the following property of the delta function:
+$\delta(f(x))=\sum_{i}\frac{1}{|\mathrm{d}f(a_{i})/\mathrm{d}x|}\delta(x-a_{i}),$ (S6)
+where $a_{i}$ is the zeros of the function $f(x)$, we can carry out the integration over $p$ and $k_{y}$ and arrive at
+$R(r)=\Theta(4-r)\frac{2}{\pi}\frac{1}{\sqrt{16-r^{2}}}.$ (S7)
+Here, $\Theta(r)$ is the step function. Finally, we obtain the effective DOS for the square lattice as
+$D(r)=\frac{\Theta(4-r)}{\pi^{2}}\frac{1}{r\sqrt{16-r^{2}}}.$ (S8)
+Appendix B Detailed calculation of the order parameter and the condensation energy for the constant effective density of states
+We show the detailed calculation of the order parameter and the condensation energy for the constant effective DOS given in Eq.
+(11) in the main text.
+First, we show the analytical solution of the NH gap equation (10) in the main text by taking into account the fact that the physical order parameter should be real.
+We can rewrite the NH gap equation for the constant effective DOS as
+$\frac{1}{U}$ $=\frac{1}{\pi\epsilon_{R}^{2}}\int\mathrm{d}\text{Re}\epsilon\mathrm{d}\text{Im}\epsilon\frac{1}{2\sqrt{\epsilon^{2}+\Delta_{0}^{2}}}$
+$=\frac{1}{\pi\epsilon_{R}^{2}}\int_{0}^{\epsilon_{R}}\mathrm{d}r\int_{-\pi}^{\pi}\mathrm{d}\theta\frac{r}{2\sqrt{r^{2}e^{2i\theta}+\Delta_{0}^{2}}}$
+$=\frac{1}{\pi\epsilon_{R}^{2}}\int_{0}^{\epsilon_{R}}\mathrm{d}r2\pi rR_{1}(r),$ (S9)
+where
+$R_{1}(r)$ $=\frac{1}{4\pi r}\int_{-\pi}^{\pi}\mathrm{d}\theta\frac{r}{\sqrt{r^{2}e^{2i\theta}+\Delta_{0}^{2}}}$
+$=-i\frac{1}{4\pi r}\oint_{|z|=1}\mathrm{d}z\frac{1}{z\sqrt{z^{2}+(\Delta_{0}/r)^{2}}},$ (S10)
+and we have introduced $z=e^{i\theta}$. By performing the contour integral with the use of the residue theorem, we get
+$R_{1}(r)$ $=\begin{cases}\frac{1}{2\Delta_{0}},&\quad(\Delta_{0}>r),\\ \frac{1}{2\Delta_{0}}-\frac{1}{\pi\Delta_{0}}\arctan(\frac{\sqrt{r^{2}-\Delta_{0}^{2}}}{\Delta_{0}}),&\quad(\Delta_{0}\leq r).\end{cases}$ (S11)
+We note that this calculation is also applicable for the square and cubic lattices with $\delta=1$.
+Here, we have to pay attention to the fact that the pole of the integrand lies inside the contour for $\Delta_{0}\leq r$.
+Then, by perfoming the integration in Eq.
+(S9) over $r$, we obtain the order parameter given in Eq.
+(12) in the main text.
+Similarly, we can analytically evaluate the condensation energy for the constant effective DOS as
+$E_{\text{cond}}$ $=\frac{\Delta_{0}^{2}}{U}-\frac{1}{\pi\epsilon_{R}^{2}}\int\mathrm{d}\text{Re}\epsilon\mathrm{d}\text{Im}\epsilon(\sqrt{\epsilon^{2}+\Delta_{0}^{2}}-\sqrt{\epsilon^{2}})$
+$=\frac{\Delta_{0}^{2}}{U}-\frac{1}{\pi\epsilon_{R}^{2}}\int_{0}^{\epsilon_{R}}\mathrm{d}r2\pi rR_{2}(r),$ (S12)
+where
+$R_{2}(r)$ $=\frac{1}{2\pi r}\int_{-\pi}^{\pi}\mathrm{d}\theta(r\sqrt{r^{2}e^{2i\theta}+\Delta_{0}^{2}}-r^{2}\sqrt{e^{2i\theta}}).$ (S13)
+We proceed with the calculation by using the contour integration, obtaining
+[…]
+Then, perfoming the integral over $r$, we get Eq.
+(13) in the main text.
+We note that we have introduced the dimensionless parameters in the main text as $E_{\text{cond}}^{\prime}\equiv E_{\text{cond}}\pi/\epsilon_{R}$, $U^{\prime}\equiv U/(\pi\epsilon_{R})$, and $\Delta^{\prime}\equiv\Delta_{0}/\epsilon_{R}$.
+For $\Delta^{\prime}\ll 1$, the condensation energy behaves asymptotically as $E_{\text{cond}}^{\prime}\sim-\Delta^{\prime 4}/6$, which is different from the results in the NH system for the constant DOS, where the condensation energy behaves as $E_{\text{cond}}^{\prime}\sim-\Delta^{\prime 2}$ for weak dissipation [93].
+Appendix C Evaluation of the order parameter at $U\to 0$ on the square lattice for $\gamma=1$
+We analytically evaluate the order parameter, which is numerically obtained in Fig.
+6 in the main text.
+Though it is difficult to get the full analytic form of the order parameter, we can estimate the exponential decay at $U\to 0$ limit for the square lattice by using
+$D(r)=\begin{cases}1/(4\pi^{2}r),&\;(r\leq 2\pi),\\ 0,&\;(r>2\pi),\end{cases}$ (S15)
+which reflects the singularity of the effective DOS at $|\epsilon|=0$.
+By using the fact that Eq.
+(S11) is applicable to this case and performing the integration over $r$, the NH gap equation is rewritten as
+[…]
+For $\Delta_{0}\ll 1$, we get
+$\frac{1}{U}\sim-\frac{1}{2\pi^{2}}\log\Delta_{0}.$ (S17)
+Then, we obtain $\Delta_{0}=\exp(-2\pi^{2}/U)$, which correctly captures the exponential decay at $U\to 0$. This indicates that the order parameter becomes finite for arbitrarily small $U$.